OpenStudy (mayliz):

what is x in x*x+12x-27

OpenStudy (bahrom7893):

what is it equal to? 0?

OpenStudy (mayliz):

=0

OpenStudy (bahrom7893):

x^2+12x-27=0 D=144+108=252

OpenStudy (bahrom7893):

x1 = (-12-sqrt(252))/2 x2 = (-12+sqrt(252))/2

OpenStudy (mayliz):

thank you